40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 9 17:12:52 2020, gautam, Update, BHD, In-air BHD OMCchamber.pdfAP_Table_20180328.pdf
    Reply  Tue Jul 14 15:29:25 2020, gautam, Update, BHD, In-air BHD - DCPD amplifier noise DCPDamp.pdf
    Reply  Wed Jul 15 19:11:40 2020, gautam, Update, BHD, In-air BHD - alignment into OMC 
       Reply  Thu Jul 16 01:12:22 2020, gautam, Update, BHD, In-air BHD - preparing the LO path IMG_8626.JPGIMG_8627.JPGIMG_8628.JPGIMG_8629.JPG
          Reply  Sun Jul 19 15:40:15 2020, gautam, Update, BHD, In-air BHD - CDS and wiring summary wiringDiagram.pdf
             Reply  Wed Jul 29 11:57:59 2020, rana, Update, BHD, In-air BHD - CDS and wiring summary 
          Reply  Mon Jul 20 17:55:15 2020, gautam, Update, BHD, In-air BHD - preparing the LO path 
             Reply  Tue Jul 21 00:30:24 2020, gautam, Update, BHD, In-air BHD - LO RIN LO_RIN.pdf
Message ID: 15489     Entry time: Thu Jul 16 01:12:22 2020     In reply to: 15483     Reply to this: 15493   15495
Author: gautam 
Type: Update 
Category: BHD 
Subject: In-air BHD - preparing the LO path 

Attachment #1 - The 80mW pickoff was getting clipped on a BNC cable, and not making it to the doubling oven. 😢 .

  • Since the PSL doubled beam isn't used for locking these days, I just didn't notice.
  • I blame the ringdown team, this crazy tee arrangement wasn't the case before.
  • I fixed the situation by changing the cabling such that the beam clears the cables comfortably.

Attachment #2 - PSL green shutter removed. Alignment into the doubling oven is extremely tedious, and so I opted to preserve the capability of recovering the green beam by simply removing a single mirror.

Attachment #3 - The beam path for coupling the LO beam into a fiber.

  • Primary goal was to have easy access to some steering mirrors so that I can optimize alignment into the fiber collimator.
  • I opted to use the NW corner of the PSL table - that's where most of our existing fiber hardware is anyways, and there was sufficient space and easy access over there.
  • 3 Y1 mirrors were installed, using the preferred Polaris mounts and 3/4" post + baseplate hardware. They were labelled Y1-1037-45P so that future workers need not be un-necessarily tortured. The third mirror is not visible in this photograph.
  • Once the collimator arrives, I will mode match this beam into the fiber. Plan is to use the fiber originally used for the mode spectroscopy project. It needs to be moved to the NW corner of the PSL table, and the other end needs to be routed to the AP table (it was brought back to the PSL table to facilitate Anjali's fiber MZ experiment). 
  • There is plenty of space in the beam path for mode-matching lens(es) and polarization control optics.

Attachment #4 shows the BHD photodiodes taken from QIL. 

  • Unfortunately, we could not find the readout electronics. 
  • In the worst case, we can just interface these PDs with the existing Satellite box (associated with the copper OMC).
  • It might be that the OMC cavity can simply be placed on this breadboard, making the whole setup nice and portable.
  • We may want to consider having an OFI between the OMC and the IFO AS beam at some point...
Attachment 1: IMG_8626.JPG  3.050 MB  Uploaded Thu Jul 16 02:12:28 2020  | Hide | Hide all
IMG_8626.JPG
Attachment 2: IMG_8627.JPG  3.513 MB  Uploaded Thu Jul 16 02:12:43 2020  | Hide | Hide all
IMG_8627.JPG
Attachment 3: IMG_8628.JPG  2.904 MB  Uploaded Thu Jul 16 02:12:54 2020  | Hide | Hide all
IMG_8628.JPG
Attachment 4: IMG_8629.JPG  2.867 MB  Uploaded Thu Jul 16 02:23:14 2020  | Hide | Hide all
IMG_8629.JPG
ELOG V3.1.3-